OpenStudy (anonymous):

if the velocity is sqrt 180 - 16x find acceleration?

OpenStudy (lgbasallote):

acceleration is the derivative of velocity

OpenStudy (lgbasallote):

so what's \[\frac{d}{dx} (\sqrt{180 - 16x})\]

OpenStudy (anonymous):

that confuses

OpenStudy (lgbasallote):

hmmm..really?

OpenStudy (lgbasallote):

which part?

OpenStudy (anonymous):

let me try

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

not getting

OpenStudy (lgbasallote):

hmm you know the chain rule right?

OpenStudy (anonymous):

make sure you have the right equation. is this it? \[ \sqrt{ 180} - 16x\]

OpenStudy (lgbasallote):

\[\Large \frac{d}{dx} (\sqrt {f(x)}) \implies \frac{d}{dx} [f(x)]^{1/2} \implies \frac 12 [f(x)]^{-1/2}] \times f'(x)\]

OpenStudy (lgbasallote):

that's the chain rule for square roots

OpenStudy (anonymous):

\[ \sqrt{180 - 16x}=(180-16x)^{\frac{1}{2}}\]

OpenStudy (anonymous):

Just take the derivative of the right hand side

OpenStudy (lgbasallote):

that's power rule + chain rule

OpenStudy (anonymous):

can you just show them the derivative and let them simplify accordingly?

OpenStudy (lgbasallote):

lol @yahoo! is gone =))

OpenStudy (anonymous):

i am back ..... sorry my internet was in a problem

mathslover (mathslover):

what is x here ?

OpenStudy (lgbasallote):

so do you have an idea how to answer this now?

OpenStudy (anonymous):

chains rule

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

a= dv/dx . ax/dt

OpenStudy (anonymous):

this is a chain rule

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

a= dv/dx . dx/dt

OpenStudy (anonymous):

dx/dt = v

OpenStudy (anonymous):

Plzz help

OpenStudy (anonymous):